### v7

# Efficient Graph Optimization via Distance-Aware Graph Representation Learning

Abstract
We propose DRTR, an efficient framework that integrates distance-aware multi-hop message passing with dynamic topology refinement. Unlike standard GNNs that rely on shallow, fixed-hop aggregation, DRTR leverages both static preprocessing and dynamic resampling to capture deeper structural dependencies. A Distance Recomputator prunes semantically weak edges using adaptive attention, while a Topology Reconstructor establishes latent connections among distant but relevant nodes. This joint mechanism enables more expressive and robust graph representation optimization across evolving graph structures. Extensive experiments demonstrate that DRTR outperforms baseline GNNs in both accuracy and scalability, with at most 20% computational overhead, especially in complex and noisy graph environments.
1 Introduction
Graph Neural Networks (GNNs) have emerged as a powerful paradigm for semi-supervised node classification and graph representation learning, particularly due to their ability to propagate information over graph structures via message passing kipf2016semi; hamilton2017inductive. Classic GNNs operate by aggregating features from multi-hop neighborhoods using static architectures and fixed aggregation depths. However, such designs may fall short when handling graphs with noisy connections, uneven structural densities, or dynamically evolving topologies Liu2021graphres; wang2021esti, leading to specific failure modes such as over-smoothing in deep layers, noise over-amplification through multi-hop propagation, and hub-dominated aggregation that biases representations toward high-degree nodes.
Recent studies have started exploring adaptive multi-hop aggregation. For example, GAMLP Zhang_2022 applies distinct MLPs for each hop-wise representation, while ImprovingTE Yao2023ImprovingTE enriches hop-specific encodings with contextual substructures. Nevertheless, these methods still depend heavily on fixed-hop sampling strategies, which may hinder generalization and limit the ability to capture complex or noisy graph signals.
In this work, we view graph optimization from a representation-centric perspective: many downstream optimization tasks—such as routing, matching, recommendation, and property prediction—critically rely on the quality of node and edge representations. DRTR optimizes the computational graph by refining multi-hop neighborhoods and reconstructing topology, thereby serving as an efficient optimization layer that improves the solution quality of downstream graph optimization problems. We propose a novel adaptive reconstruction framework for multi-hop GNNs that dynamically adjusts both distance computations and local graph structures to promote more effective and robust message passing. Inspired by coreset selection guo2022deepcore, which selects a small, informative subset of data points to approximate the full dataset, our method adaptively reconstructs the neighborhood for each node by selecting a small, informative subset of neighbors to approximate the full multi-hop neighborhood, enabling the model to better capture latent semantic patterns and mitigate the propagation of noise.
Specifically, we introduce two core modules: the Distance Recomputator and the Topology Reconstructor, which together form our DRTR framework. The Distance Recomputator recalibrates node proximity metrics based on adaptive structural signals, refining how multi-hop neighbors are selected and weighted during message passing. Meanwhile, the Topology Reconstructor adjusts local edges to reflect more meaningful structural patterns, thereby enhancing the resilience of the model to noisy or adversarial graph changes.
Unlike previous models that statically aggregate neighborhood features, DRTR continuously adapts its perception of graph topology and node relationships throughout training. This dynamic reconstruction mechanism improves the quality of learned representations, stabilizes optimization dynamics, and enables better generalization across a wide range of graphs.
To validate the effectiveness of our framework, we conduct comprehensive experiments across several benchmark datasets, including both static and dynamic graph settings. Our results demonstrate that DRTR consistently outperforms state-of-the-art multi-hop GNN models in terms of accuracy, robustness, and stability. Furthermore, our theoretical analysis sheds light on why adaptive graph refinement enhances graph representation optimization by aligning information flow with the underlying graph semantics.
Our contributions are threefold:
(i) We introduce DRTR, a distance-aware multi-hop framework that explicitly separates diffusion, distance recomputation, and topology reconstruction, enabling robust aggregation under noisy and evolving graphs.
(ii) We design two complementary modules—the Distance Recomputator (DR) and Topology Reconstructor (TR)—which jointly prune noisy edges and add latent long-range links in a task-driven manner, improving representation quality and model robustness in both static and evolving graph scenarios.
(iii) We provide empirical and theoretical analysis showing that DRTR reduces effective neighborhood size and improves generalization: empirically, DRTR consistently improves accuracy and stability across diverse benchmarks and downstream tasks; theoretically, we establish generalization, convergence, and stability guarantees under adaptive topology refinement.
2 Related Works
Unlike classical combinational graph optimization (e.g., shortest paths, routing, matching), representation-centric graph optimization focuses on optimizing the message-passing pathways and refining graph structure to support downstream optimization tasks. DRTR belongs to this second class, where graph optimization refers to optimizing graph structure, optimizing message passing paths, and improving downstream optimization tasks.
Adaptive Structure Learning in GNNs
Conventional Graph Neural Networks (GNNs) such as GCN kipf2016semi, GraphSAGE hamilton2017inductive, and GAT vel2018graph adopt fixed-hop message passing schemes and pre-defined neighbor aggregation strategies. However, recent studies recognize that structural connections in graphs often suffer from noise, imbalance, or redundancy liu2023survey. To address this, various works have explored structure learning franceschi2019learning; chen2020iterative; Liu2021graphres, which attempts to refine or reconstruct the graph topology during training. Nonetheless, many such methods rely on end-to-end optimization or learn a static edge mask, which lacks the flexibility to respond to evolving node semantics during multi-hop propagation.
Distance-aware Message Passing
Several studies have attempted to incorporate distance or importance-based filtering into message passing. For instance, Personalized PageRank (PPR)-based methods gasteiger2018pp; chien2021adaptive reweight neighbor contributions based on precomputed transition probabilities. Meanwhile, GAMLP Zhang_2022 and ImprovingTE Yao2023ImprovingTE explicitly leverage multi-hop propagation, with hop-specific encoders and contextual subgraph tokens. However, these approaches still depend on fixed-hop sampling, which may not fully capture varying node proximities or information decay across hops. Unlike prior work, DRTR dynamically recomputes node distances during training using the proposed Distance Recomputator, allowing for more context-aware and adaptive neighborhood construction.
Asynchronous and Selective Aggregation
Traditional GNNs typically perform synchronous aggregation, where all neighbor messages are processed simultaneously within each hop wu2020comprehensive. This design overlooks the fact that different neighbors may carry drastically different relevance to the target node. Asynchronous strategies li2017gated; 9878467 attempt to stagger updates or introduce temporal dynamics, but often target streaming or event-based graphs. In contrast, our approach introduces a selective and relevance-aware aggregation mechanism by jointly optimizing node distance and topology. The Topology Reconstructor module adaptively rewires the local structure, filtering out low-utility links and amplifying informative ones across multiple hops.
Heat Diffusion in Graph Representation Learning
The concept of heat diffusion has inspired a line of works modeling how signals attenuate over graphs thanou2016learning; Chen_2021_prop; xu2020heatk. These models typically employ analytical kernels to encode diffusion patterns into the graph structure. However, they are often static and not integrated with task-driven learning. Our proposed DRTR framework incorporates a diffusion-inspired attenuation mechanism, which adaptively controls information decay based on recomputed distances, enabling fine-grained signal control and improved robustness against over-smoothing or noise amplification in deep GNNs.
In summary, DRTR bridges the gap between dynamic topology learning, relevance-aware aggregation, and distance-sensitive diffusion, providing a unified framework that extends beyond existing adaptive GNN approaches.
3 Methodology
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ denote an undirected graph with node set $\mathcal{V}$ and edge set $\mathcal{E}$. Each node $v\in\mathcal{V}$ has an associated input feature $\mathbf{x}_{v}\in\mathbb{R}^{d}$. Let $\mathbf{A}\in\mathbb{R}^{N\times N}$ be the adjacency matrix and $\mathbf{X}\in\mathbb{R}^{N\times d}$ the feature matrix. Our goal is to predict node labels $y_{v}$ for $v\in\mathcal{V}_{\text{unlabeled}}$ using a subset of labeled nodes $\mathcal{V}_{L}$.
Unlike standard GNNs that perform shallow 1-hop message passing layer-by-layer, DRTR explicitly models $K$-hop diffusion via direct aggregation across multi-hop neighborhoods. To prevent topological noise propagation, DRTR introduces two key modules: the Distance Recomputator, which recalibrates semantic distances between nodes in the multi-hop domain and prunes noisy edges, and the Topology Reconstructor, which establishes latent high-quality connections between semantically related but structurally distant nodes.
3.1 Multi-Hop Diffusion Aggregation
Formally, we denote $\mathcal{N}^{(k)}(v)$ as the $k$-hop neighborhood of node $v$. Instead of stacking $K$ layers of 1-hop aggregation, we directly aggregate from each $k$-hop neighborhood via a heat-inspired attention mechanism:
$\mathbf{h}_{v}^{(k)}=\sum_{u\in\mathcal{N}^{(k)}(v)}\alpha_{vu}^{(k)}\cdot\mathbf{W}^{(k)}\mathbf{x}_{u},$ (1)
where the attention coefficient $\alpha_{vu}^{(k)}$ is defined as:
$\alpha_{vu}^{(k)}=\frac{\exp\left(\text{LeakyReLU}(\mathbf{a}^{\top}[\mathbf{W}^{(k)}\mathbf{x}_{v}\,\|\,\mathbf{W}^{(k)}\mathbf{x}_{u}])/\tau_{k}\right)}{\sum_{u^{\prime}\in\mathcal{N}^{(k)}(v)}\exp\left(\text{LeakyReLU}(\mathbf{a}^{\top}[\mathbf{W}^{(k)}\mathbf{x}_{v}\,\|\,\mathbf{W}^{(k)}\mathbf{x}_{u^{\prime}}])/\tau_{k}\right)}.$ (2)
Here, $\mathbf{W}^{(k)}$ is a hop-specific transformation matrix that enables different feature transformations at different hop distances, while $\mathbf{a}$ is a shared attention vector.
The temperature parameter $\tau_{k}$ controls the sharpness of attention distribution and follows a decay schedule:
$\tau_{k}=\tau_{0}\cdot\exp(-\eta k),\quad\eta>0,$ (3)
where $\tau_{0}$ is the initial temperature and $\eta$ controls the decay rate across hops.
To ensure numerical stability and prevent vanishing gradients, we apply layer normalization after each hop aggregation:
$\tilde{\mathbf{h}}_{v}^{(k)}=\text{LayerNorm}\left(\mathbf{h}_{v}^{(k)}\right),$ (4)
which normalizes the activations to have zero mean and unit variance (see Appendix for detailed formulation).
The final representation for node $v$ is then given by:
$\mathbf{z}_{v}=\textsf{COMBINE}\left(\{\tilde{\mathbf{h}}_{v}^{(k)}\}_{k=1}^{K}\right)=\sum_{k=1}^{K}\gamma_{k}\cdot\tilde{\mathbf{h}}_{v}^{(k)},\quad\sum_{k}\gamma_{k}=1.$ (5)
The combination weights $\gamma_{k}$ are learned through a softmax parameterization:
$\gamma_{k}=\frac{\exp(\phi_{k})}{\sum_{j=1}^{K}\exp(\phi_{j})},$ (6)
where $\{\phi_{k}\}_{k=1}^{K}$ are learnable parameters that control the relative importance of different hop distances. We use global hop-level weights $\gamma_{k}$ rather than per-node gating (e.g., $\gamma_{k}(v)=\text{softmax}(\bm{\psi}_{k}^{\top}\tilde{\mathbf{h}}_{v}^{(k)})$) to maintain computational efficiency and ensure consistent hop importance across the graph, which is particularly important for distance-aware aggregation where hop semantics should be globally consistent.
3.2 Distance Recomputator (DR)
This module performs local graph optimization by recalibrating node-to-node distances and removing edges that degrade downstream optimization quality. To mitigate topological noise, we introduce a dynamic Distance Recomputator module that filters weak edges based on learned semantic distance. The pairwise distance between $v$ and $u\in\mathcal{N}^{(k)}(v)$ is defined as:
$d_{vu}^{(k)}=\left\|\mathbf{x}_{v}-\mathbf{x}_{u}\right\|_{2}^{2}+\lambda_{k}\cdot\delta^{(k)}_{vu}$, (7)
where $\delta^{(k)}_{vu}$ is a learned penalty term from diffusion depth $k$.
The penalty term $\delta^{(k)}_{vu}$ incorporates both structural and semantic information:
$\delta^{(k)}_{vu}=\beta_{1}\cdot\text{hop\_penalty}(k)+\beta_{2}\cdot\text{semantic\_divergence}(\mathbf{x}_{v},\mathbf{x}_{u}),$ (8)
where $\text{hop\_penalty}(k)=k^{2}$ penalizes longer paths, and $\text{semantic\_divergence}(\mathbf{x}_{v},\mathbf{x}_{u})=1-\cos(\mathbf{x}_{v},\mathbf{x}_{u})$ measures feature dissimilarity. The distance definition combines Euclidean distance $\|\mathbf{x}_{v}-\mathbf{x}_{u}\|_{2}^{2}$ (capturing magnitude differences) with cosine-based semantic divergence (capturing directional differences), where $\beta_{1}$ and $\beta_{2}$ control the trade-off between structural depth penalty and semantic similarity.
The scaling parameter $\lambda_{k}$ follows an adaptive schedule:
$\lambda_{k}=\lambda_{0}\cdot\exp(-\rho k)+\lambda_{\min},$ (9)
where $\lambda_{0}$ is the initial scaling, $\rho$ controls the decay rate, and $\lambda_{\min}$ ensures minimum regularization.
For each hop $k$, we discard high-distance neighbors using a soft thresholding mechanism:
$\mathcal{N}^{(k)}(v)\leftarrow\left\{u\in\mathcal{N}^{(k)}(v)\mid d_{vu}^{(k)}\leq\alpha_{k}\right\},$ (10)
where the threshold $\alpha_{k}$ is set as the $p$-th percentile of $\{d_{vu}^{(k)}\}_{u\in\mathcal{N}^{(k)}(v)}$, which adaptively controls the pruning ratio. This percentile-based approach ensures that a consistent fraction of neighbors is retained across different nodes and hops, maintaining stable aggregation patterns.
The distance recomputation process is end-to-end differentiable, enabling gradient-based optimization of the distance metric alongside the main classification objective (see Appendix for gradient details).
3.3 Topology Reconstructor (TR)
This module performs global topology optimization by adding latent edges that improve connectivity and information flow for downstream optimization tasks. The Topology Reconstructor augments the graph with new edges to improve structural balance. We identify semantically similar but topologically distant nodes based on a multi-criteria similarity function. To maintain consistency (where larger values indicate greater similarity), we define:
$\text{sim}_{vu}=\omega_{1}\cdot\text{contextual\_alignment}(v,u)+\omega_{2}\cdot\text{structural\_similarity}(v,u)-\omega_{3}\cdot\left\|\mathbf{x}_{v}-\mathbf{x}_{u}\right\|_{2}^{2},$ (11)
where $\omega_{1},\omega_{2},\omega_{3}\geq 0$ are learnable weights.
The structural similarity component captures local graph structure:
$\text{structural\_similarity}(v,u)=\frac{|\mathcal{N}(v)\cap\mathcal{N}(u)|}{|\mathcal{N}(v)\cup\mathcal{N}(u)|},$ (12)
where $\mathcal{N}(v)$ denotes the 1-hop neighborhood of node $v$.
The contextual alignment measures feature compatibility in the embedding space:
$\text{contextual\_alignment}(v,u)=\frac{\mathbf{x}_{v}^{T}\mathbf{x}_{u}}{\|\mathbf{x}_{v}\|_{2}\|\mathbf{x}_{u}\|_{2}}.$ (13)
The edge addition process follows a probabilistic approach to prevent over-connection:
[…]
where $\sigma(\cdot)$ is the sigmoid function, $\beta$ is the similarity threshold, and $\tau$ is a temperature parameter controlling the sharpness of the probability distribution.
[…]
$\mathcal{E}\leftarrow\mathcal{E}\cup\{(v,u)\},$ (15)
where $\theta$ is a probability threshold controlling edge addition.
To ensure computational efficiency, we restrict candidate edge selection to approximate $k$-nearest neighbors in the embedding space, limiting the candidate set $\mathcal{E}_{\text{candidate}}$ to at most $R$ edges per node. This reduces the complexity from $\mathcal{O}(|\mathcal{V}|^{2})$ to $\mathcal{O}(|\mathcal{V}|\cdot k\cdot d)$ where $k$ is the number of approximate nearest neighbors and $d$ is the feature dimension.
The topology reconstruction process is optimized through a margin-based objective that encourages high-similarity pairs to be connected:
$\mathcal{L}_{\text{TR}}=\sum_{(v,u)\in\mathcal{E}_{\text{candidate}}}\max(0,\beta-\text{sim}_{vu})\cdot P(\text{add edge }(v,u)),$ (16)
which penalizes high-similarity pairs that are not connected, encouraging the model to add edges between semantically similar nodes.
This process improves neighborhood expressivity and alleviates under-connection problems commonly found in real-world graphs while maintaining computational efficiency through selective edge addition.
3.4 DRTR Training Procedure
[…]
Output: Node embeddings $\mathbf{Z}\in\mathbb{R}^{N\times d^{\prime}}$
Initialize $Q_{h}^{(k)}(v)\leftarrow 0$, $N_{h}^{(k)}(v)\leftarrow 0$ for all $(v,k,h)\in\mathcal{V}\times[K]\times[K]$;
Initialize $\mathbf{W}^{(k)}\in\mathbb{R}^{d\times d^{\prime}}$, $\phi_{k}$ for all $k\in[K]$, $\gamma_{k}\leftarrow\text{softmax}(\phi_{k})$;
for training iteration $t=1,\ldots,T$ do
 for each node $v\in\mathcal{V}$ do
    for $k=1,\ldots,K$ do
       $\mathcal{N}^{(k)}(v)\leftarrow\text{StaticSampling}(\mathcal{G},v,k)$;
         Compute $\alpha_{vu}^{(k)}$ via Eq. (2);
       $\mathbf{h}_{v}^{(k)}\leftarrow\sum_{u\in\mathcal{N}^{(k)}(v)}\alpha_{vu}^{(k)}\cdot\mathbf{W}^{(k)}\mathbf{x}_{u}$;
       $\tilde{\mathbf{h}}_{v}^{(k)}\leftarrow\text{LayerNorm}(\mathbf{h}_{v}^{(k)})$;
      end for
    $\mathbf{z}_{v}\leftarrow\sum_{k=1}^{K}\gamma_{k}\cdot\tilde{\mathbf{h}}_{v}^{(k)}$;
   end for
 for each node $v\in\mathcal{V}$ do
    for $k=1,\ldots,K$ do
       for each $u\in\mathcal{N}^{(k)}(v)$ do
          $n\leftarrow N_{k}(v,u)\leftarrow N_{k}(v,u)+1$;
          $d_{vu}^{(k)}\leftarrow\|\mathbf{x}_{v}-\mathbf{x}_{u}\|_{2}^{2}+\lambda_{k}\cdot\delta_{vu}^{(k)}$;
         end for
       $\alpha_{k}\leftarrow p\text{-th percentile of }\{d_{vu}^{(k)}\}_{u\in\mathcal{N}^{(k)}(v)}$;
       $\mathcal{N}^{(k)}(v)\leftarrow\{u\in\mathcal{N}^{(k)}(v)\mid d_{vu}^{(k)}\leq\alpha_{k}\}$;
      end for
   end for
 $\mathcal{E}_{\text{candidate}}\leftarrow\text{ApproximateKNN}(\mathbf{X},k)$;
 for each $(v,u)\in\mathcal{E}_{\text{candidate}}\setminus\mathcal{E}$ do
    $\text{sim}_{vu}\leftarrow\omega_{1}\cdot\text{contextual\_alignment}(v,u)+\omega_{2}\cdot\text{structural\_similarity}(v,u)-\omega_{3}\cdot\|\mathbf{x}_{v}-\mathbf{x}_{u}\|_{2}^{2}$;
    if $\sigma((\text{sim}_{vu}-\beta)/\tau)>\theta$ then
       $\mathcal{E}\leftarrow\mathcal{E}\cup\{(v,u)\}$;
      end if
   end for
  Compute loss $\mathcal{L}$ and update parameters via gradient descent;
end for
Algorithm 1 DRTR Training Algorithm
4 Theoretical Analysis
In this section, we provide a theoretical foundation for the DRTR framework by analyzing how adaptive topology reconstruction enhances generalization in graph learning. We focus on the interaction between semantic distance calibration, dynamic edge pruning/addition, and the impact on information propagation under the message passing paradigm.
4.1 Preliminaries
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be an undirected graph, and let $\mathbf{X}\in\mathbb{R}^{N\times d}$ be the input feature matrix. The DRTR framework constructs multi-hop diffusion representations via:
$\mathbf{z}_{v}=\sum_{k=1}^{K}\gamma_{k}\cdot\mathbf{h}_{v}^{(k)}=\sum_{k=1}^{K}\gamma_{k}\cdot\sum_{u\in\mathcal{N}^{(k)}(v)}\alpha_{vu}^{(k)}\cdot\mathbf{W}^{(k)}\mathbf{x}_{u}.$
Here, $\alpha_{vu}^{(k)}$ encodes attention over $k$-hop neighbors, and $\mathcal{N}^{(k)}(v)$ is dynamically adjusted by the Distance Recomputator (DR) and Topology Reconstructor (TR).
We aim to study the generalization performance of DRTR under a semi-supervised node classification setting, where the loss is defined over a labeled subset $\mathcal{V}_{L}\subset\mathcal{V}$.
4.2 Main Theoretical Results
The theoretical results show that DRTR optimizes the computational graph by reducing noise propagation and improving connectivity, which are core primitives underlying many graph optimization problems. By tightening generalization bounds and improving stability under perturbations, DRTR provides theoretical guarantees for efficient graph optimization in large, noisy, or evolving graphs.
Our main theoretical contribution focuses on how DRTR’s adaptive neighborhood pruning reduces effective degree and improves generalization:
Theorem 4.1 (Generalization Bound under Adaptive Neighborhood Pruning).
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph with $|\mathcal{V}|=n$ nodes, and let $d_{\text{eff}}$ denote the average effective degree after DRTR’s distance-based pruning. Assume the underlying ground-truth graph satisfies structural smoothness, and DRTR correctly removes $\epsilon$-fraction noisy edges and adds $\eta$-fraction semantically valid edges. Then, with probability at least $1-\delta$:
$\mathcal{L}_{\text{true}}\leq\mathcal{L}_{\text{emp}}+\mathcal{O}\left(\sqrt{\frac{d_{\text{eff}}\cdot\log n}{|\mathcal{V}_{L}|}}+\sqrt{\frac{\log(1/\delta)}{|\mathcal{V}_{L}|}}\right),$
where $d_{\text{eff}}=\frac{1}{n}\sum_{v\in\mathcal{V}}|\mathcal{N}_{\text{eff}}(v)|$ and $\mathcal{N}_{\text{eff}}(v)=\bigcup_{k=1}^{K}\{u\in\mathcal{N}^{(k)}(v):d_{vu}^{(k)}\leq\alpha_{k}\}$ is the effective neighborhood after pruning. The key insight is that $d_{\text{eff}}\ll d_{\text{original}}$ due to distance-based filtering, leading to reduced hypothesis complexity and improved generalization.
The detailed proof, which establishes the connection between effective degree reduction and Rademacher complexity bounds, is provided in the Appendix.
Proposition 4.2 (Convergence Rate).
Under standard assumptions (bounded gradients, Lipschitz continuity), and noting that DRTR’s distance recomputation and topology reconstruction modules preserve $L$-smoothness and bounded gradient assumptions, the DRTR algorithm converges to a stationary point at rate $\mathcal{O}(1/\sqrt{T})$, where $T$ is the number of iterations. The proof follows standard SGD analysis (see Appendix).
Proposition 4.3 (Stability under Graph Perturbation).
Let $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ be two graphs differing in at most $\Delta$ edges. Let $\mathbf{Z}_{1}$ and $\mathbf{Z}_{2}$ be the representations learned by DRTR on $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, respectively. Then:
$\|\mathbf{Z}_{1}-\mathbf{Z}_{2}\|_{F}\leq C\cdot\Delta\cdot\sqrt{|\mathcal{V}|},$
where $C$ is a constant depending on the Lipschitz constant and graph properties. This stability guarantee explains DRTR’s robustness to noisy and dynamically evolving graphs (see Appendix for proof).
4.3 Discussion
DRTR improves robustness and generalization by reducing variance (selective pruning), enhancing representation power (topology completion), and regularizing the objective (adaptive neighborhood smoothing). Detailed proofs are in the Appendix.
5 Experimental Analysis
We evaluate DRTR from the perspective of graph optimization quality, where optimization refers to improving the structure, connectivity, and effective neighborhoods that downstream tasks rely on. We conduct comprehensive experiments to evaluate the effectiveness of our proposed DRTR framework. We evaluate three DRTR-based variants:
  • GDRA: baseline + Distance Recomputator (DR only, Section 3.2);
  • GKHDA: baseline + K-hop heat diffusion aggregator (Section 3.1) without DR/TR;
  • GKHDDRA: full DRTR with diffusion, DR, and TR (Sections 3.1–3.3).
We address three key questions: (Q1) How does DRTR improve performance across diverse datasets and GNN backbones? (Q2) What is the contribution of each module (DR, TR, Diffusion)? (Q3) Can DRTR generalize to real-world downstream tasks such as recommendation or molecular modeling?
5.1 Datasets and Baselines
We evaluate on five benchmark datasets: Cora, Citeseer, Pubmed (standard citation graphs), ogbn-arxiv, and ogbn-products (from OGB hu2020open). Each dataset varies in terms of scale, edge density, and semantic complexity.
Baselines include standard GNNs (GCN, SGC, SSGC, GAT, GraphSAGE, APPNP) and their DRTR-augmented counterparts. We adopt 3-layer configurations and use early stopping based on validation accuracy. All results are averaged over 10 random seeds.
[…]
tab:all-results Model Cora Citeseer Pubmed ogbn-arxiv ogbn-products Acc Time(s) Var Acc Time(s) Var Acc Time(s) Var Acc Acc GCN 81.2 3.5 0.021 70.9 3.6 0.025 79.3 12.7 0.018 70.5 75.4 GCN+GDRA 82.6 3.9 0.016 71.3 3.8 0.019 80.1 12.5 0.016 73.1 76.6 GCN+GKHDA 82.4 4.0 0.017 71.7 4.1 0.018 80.5 12.4 0.016 72.8 76.1 GCN+GKHDDRA 82.7 4.2 0.013 72.3 4.3 0.014 80.9 12.3 0.014 73.9 77.2 SGC 74.2 2.1 0.030 71.5 2.3 0.026 78.2 10.4 0.024 68.2 74.1 SGC+GDRA 75.8 2.3 0.022 73.1 2.5 0.021 81.2 10.3 0.019 70.3 75.2 SGC+GKHDA 75.1 2.4 0.021 73.4 2.6 0.020 81.6 10.2 0.018 70.0 75.0 SGC+GKHDDRA 77.4 2.6 0.018 74.6 2.8 0.017 82.5 10.1 0.017 71.2 76.3 SSGC 83.0 2.5 0.015 75.6 2.6 0.017 73.6 11.3 0.016 69.9 74.8 SSGC+GDRA 83.2 2.6 0.013 76.4 2.7 0.015 74.2 11.1 0.014 71.1 75.4 SSGC+GKHDA 84.3 2.7 0.012 76.1 2.8 0.014 74.5 11.0 0.013 71.5 75.6 SSGC+GKHDDRA 84.1 2.8 0.011 77.6 2.9 0.013 74.7 10.9 0.012 72.0 76.8 APPNP 82.3 4.5 0.017 75.2 4.7 0.021 71.5 13.2 0.020 71.6 75.9 APPNP+GDRA 83.5 4.7 0.014 74.4 4.9 0.018 73.6 13.0 0.018 72.9 76.5 APPNP+GKHDA 83.8 4.9 0.013 74.5 5.0 0.016 74.1 12.9 0.016 73.1 76.3 APPNP+GKHDDRA 84.6 5.1 0.011 75.3 5.2 0.014 74.5 12.8 0.014 74.3 77.6
Table 1: Accuracy (%), convergence time (s), and variance across datasets for baseline and DRTR-augmented models. GKHDDRA consistently improves both performance and stability.
Q1: Accuracy and Stability Across Datasets
[…]
fig:dataset-comparison Refer to caption
Figure 1: Performance across datasets. (a) Accuracy heatmap showing improvements across all model-dataset combinations. (b) Average improvement by dataset. (c) Variance reduction demonstrating improved stability. (d) Best performance comparison. (e) Training time overhead showing modest computational cost.
[…]
fig:comprehensive-performance provides a comprehensive performance analysis across multiple dimensions. Panel (a) shows accuracy improvements by model, where DRTR variants consistently outperform baselines across GCN, SGC, SSGC, and APPNP. Panel (b) demonstrates variance reduction, with GKHDDRA achieving up to 38.1% variance reduction on Cora. Panel (c) illustrates training time overhead, showing that DRTR introduces only modest computational overhead (4-20% depending on model). Panel (d) presents accuracy comparison across all five datasets using GCN backbone, demonstrating consistent improvements. Panel (e) shows component contributions, confirming that the full DRTR achieves the best performance.
We observe several consistent trends. Across all datasets and models, DRTR variants outperform their corresponding baselines, with the full version (GKHDDRA) achieving the best accuracy in nearly all settings. For example, SGC improves from 74.2% to 77.4% on Cora (+3.2%), and APPNP improves from 82.3% to 84.6% (+2.3%). On average, GKHDDRA achieves 1.5–3.4% accuracy improvement across datasets, with the largest gains on ogbn-arxiv (3.4%). In all cases, DRTR-enhanced models exhibit lower performance variance across training runs, indicating more stable optimization dynamics and reduced sensitivity to initialization. For instance, variance reduction ranges from 23.8% (GDRA on Cora) to 44% (GKHDDRA on Citeseer), with an average reduction of 38.1% for the full DRTR variant.
While DRTR introduces additional computations (notably in the diffusion and reconstruction phases), the increase in training time remains modest (4–20% depending on model). For instance, GCN’s training time on Pubmed decreases slightly from 12.7s to 12.3s (with DRTR), indicating that dynamic reconstruction does not introduce noticeable overhead and can even stabilize convergence (leading to fewer epochs to early stopping). Some models show improved training time due to faster convergence to optimal performance. Improvements hold across diverse architectures, including shallow models like SGC and diffusion-style models like APPNP and SSGC, suggesting that DRTR can be used as a generic plug-in for a wide range of GNNs.
Overall, these results highlight DRTR’s effectiveness as a general-purpose graph enhancement module—improving both predictive power and training consistency across scales, backbones, and graph domains.
[…]
fig:comprehensive-performance Refer to caption
Figure 2: Comprehensive performance analysis. (a) Accuracy improvement by model showing consistent gains across GCN, SGC, SSGC, and APPNP. (b) Variance reduction demonstrating improved stability. (c) Training time overhead showing modest computational cost. (d) Accuracy across datasets using GCN backbone. (e) Component contribution analysis on Cora.
Q2: Component Analysis
[…]
Table LABEL:tab:component-mini quantifies the individual contributions: GDRA (Distance Recomputator) alone provides +1.4% accuracy improvement and 23.8% variance reduction on Cora, GKHDA (K-hop Heat Diffusion Aggregator) provides +1.2% improvement and 19.0% variance reduction, while TR contributes an additional +0.3% improvement and 18.8% variance reduction. The full DRTR achieves +1.5% improvement and 38.1% variance reduction, demonstrating that the modules work synergistically rather than additively. GDRA removes noisy links and helps stabilize learning, GKHDA enhances long-range message passing, and GKHDDRA combines the above with topology reconstruction for maximal performance.
[…]
fig:component-ablation Refer to caption
Figure 3: Component ablation analysis. (a) Accuracy improvement by component showing individual and combined contributions. (b) Variance reduction demonstrating stability improvements. (c) Combined effect analysis showing the optimal trade-off achieved by full DRTR.
[…]
tab:component-mini Module Improvement $\uparrow$ Variance $\downarrow$ GDRA +1.4% 23.8% GKHDA +1.2% 19.0% TR (in full) +0.3% 18.8% DRTR (full) +1.5% 38.1%
Table 2: Component contributions on Cora (GCN backbone).
Q3: Generalization to Real-World Downstream Tasks
To assess the broader applicability of DRTR, we extend our evaluation to two real-world downstream tasks beyond node classification. These two tasks—link prediction and molecular property prediction—are widely used as surrogate optimization benchmarks, because better graph representations lead to better optimization performance (ranking quality or regression accuracy):
Task 1: Recommendation (Link Prediction)
We apply DRTR to a recommendation setting based on the MovieLens-100K dataset, where user–item interactions are modeled as a bipartite graph. The goal is to predict missing user–item links.
We follow a standard link prediction pipeline using node embeddings from a DRTR-augmented GraphSAGE model. Evaluation metrics include AUC and Average Precision (AP):
[…]
tab:recsys Model AUC AP GraphSAGE 93.1 91.7 GraphSAGE+GDRA 94.0 92.5 GraphSAGE+GKHDA 94.3 92.8 GraphSAGE+GKHDDRA 95.1 93.6
Table 3: Link prediction results on MovieLens-100K. DRTR improves both ranking metrics.
Task 2: Molecular Property Prediction
We evaluate DRTR on the ZINC-12K dataset for molecular graph regression, using a GCN-based backbone. Each molecule is represented as a graph with atoms as nodes and bonds as edges. The task is to predict molecular properties (e.g., solubility, logP).
We report Mean Absolute Error (MAE, ↓) across three targets:
[…]
tab:zinc Model logP QED SA GCN 0.423 0.218 0.387 GCN+GDRA 0.401 0.205 0.375 GCN+GKHDA 0.395 0.203 0.372 GCN+GKHDDRA 0.383 0.197 0.366
Table 4: Molecular property prediction on ZINC-12K (MAE ↓). DRTR enhances representation in dense chemical graphs.
Summary
These results demonstrate DRTR’s ability to improve ranking and predictive performance in sparse user–item graphs (recommendation) while preserving fine-grained local-global dependencies in dense, noisy molecular graphs. On MovieLens-100K, DRTR achieves relative improvements of +2.0% AUC (from 93.1% to 95.1%) and +1.9% AP (from 91.7% to 93.6%) over the baseline GraphSAGE. On ZINC-12K, DRTR reduces MAE by 9.5% for logP (from 0.423 to 0.383), 9.6% for QED (from 0.218 to 0.197), and 5.4% for SA (from 0.387 to 0.366), demonstrating consistent improvements across different molecular properties.
This validates DRTR’s generality across structural domains—sparse bipartite graphs, dense chemical graphs, and traditional citation networks—with relative performance gains ranging from 2.0% to 9.6% depending on the task and metric.
6 Conclusion
In this paper, we propose DRTR, a distance-aware graph representation optimization framework that can optimize GNNs by dynamically refining multi-hop neighborhoods and topological structures. DRTR integrates a Distance Recomputator to prune noisy links based on semantic similarity, and a Topology Reconstructor to add meaningful long-range connections. Combined with a heat-diffusion-based message passing scheme, DRTR enables robust and expressive node representations. Extensive experiments across node classification, link prediction, and molecular property prediction confirm that DRTR consistently improves accuracy, stability, and generalization, while maintaining scalability across both synthetic and real-world graphs. Taken together, DRTR provides an efficient and general optimization layer that improves graph structure, optimizes message-passing pathways, and enhances the solution quality of downstream graph optimization problems.
Appendix A Appendix
A.1 Complete Proof of Theorem 1 (Main Text): Generalization Bound
Theorem A.1 (Complete Generalization Bound under Adaptive Neighborhood Pruning).
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph with $|\mathcal{V}|=n$ nodes, and let $d_{\text{eff}}$ denote the average effective degree after DRTR’s distance-based pruning. Let $\mathcal{G}^{*}=(\mathcal{V},\mathcal{E}^{*})$ be the ground-truth graph satisfying structural smoothness. Assume the observed graph $\mathcal{G}$ contains $\epsilon$-fraction noisy edges, DRTR correctly removes $\epsilon$-fraction noisy edges and adds $\eta$-fraction semantically valid edges, the GNN is $L$-Lipschitz with respect to input features, and node features are bounded such that $\|\mathbf{x}_{v}\|_{2}\leq B$ for all $v\in\mathcal{V}$.
Then, with probability at least $1-\delta$, the generalization error satisfies:
$\mathcal{L}_{\text{true}}\leq\mathcal{L}_{\text{emp}}+\mathcal{O}\left(\sqrt{\frac{d_{\text{eff}}\cdot\log n}{|\mathcal{V}_{L}|}}+\sqrt{\frac{\log(1/\delta)}{|\mathcal{V}_{L}|}}\right),$
where $d_{\text{eff}}=\frac{1}{n}\sum_{v\in\mathcal{V}}|\mathcal{N}_{\text{eff}}(v)|$ and $\mathcal{N}_{\text{eff}}(v)=\bigcup_{k=1}^{K}\{u\in\mathcal{N}^{(k)}(v):d_{vu}^{(k)}\leq\alpha_{k}\}$ is the effective neighborhood after pruning.
Proof A.2 (Complete Proof).
We provide a detailed proof using Rademacher complexity analysis and structural graph properties. The proof proceeds in several steps, establishing key lemmas that control error propagation through the multi-hop aggregation structure.
Notation. We denote by $I[A]$ the indicator function for event $A$. For each node $v\in\mathcal{V}$ and hop $k\in[K]$, let $t=N_{k}(v)$ denote the number of times node $v$ has been processed at hop $k$ up to the current iteration. We denote by $\mathbf{z}_{v}^{t}$ the representation of node $v$ after $t$ updates at hop $k$.
Step 1: Rademacher Complexity Analysis
Let $\mathcal{H}_{\text{DRTR}}$ denote the hypothesis class of DRTR-induced representations. The Rademacher complexity is:
$\mathcal{R}_{n}(\mathcal{H}_{\text{DRTR}})=\mathbb{E}_{\sigma}\left[\sup_{h\in\mathcal{H}_{\text{DRTR}}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}h(\mathbf{x}_{i})\right],$ (17)
where $\sigma_{i}\in\{-1,+1\}$ are independent Rademacher random variables.
Step 2: Effective Neighborhood Analysis
After DRTR processing, the effective neighborhood size is reduced. Let $d_{\text{eff}}$ denote the average effective degree. For each node $v$, the number of neighbors used in aggregation is:
$|\mathcal{N}_{\text{eff}}(v)|=\sum_{k=1}^{K}|\mathcal{N}^{(k)}(v)\cap\{u:d_{vu}^{(k)}\leq\alpha_{k}\}|.$ (18)
The key insight is that $d_{\text{eff}}\ll d_{\text{original}}$ due to distance-based pruning. Specifically, by the percentile-based thresholding mechanism, we have:
$d_{\text{eff}}=\frac{1}{n}\sum_{v\in\mathcal{V}}|\mathcal{N}_{\text{eff}}(v)|\leq(1-p)\cdot d_{\text{original}},$ (19)
where $p$ is the pruning percentile parameter.
Step 3: Lipschitz Constant Analysis
The DRTR representation function $f_{\text{DRTR}}$ satisfies:
$\|f_{\text{DRTR}}(\mathbf{x}_{v})-f_{\text{DRTR}}(\mathbf{x}_{u})\|_{2}\leq L_{\text{DRTR}}\|\mathbf{x}_{v}-\mathbf{x}_{u}\|_{2},$ (20)
where $L_{\text{DRTR}}=L\cdot\prod_{k=1}^{K}(1+\lambda_{k})$ accounts for the multi-hop aggregation and distance recomputation. The Lipschitz constant can be bounded as:
$L_{\text{DRTR}}\leq L\cdot\left(1+\frac{\lambda_{0}}{H}\right)^{K}\leq L\cdot e^{\lambda_{0}K/H}=\mathcal{O}(L),$ (21)
since $\lambda_{k}=\lambda_{0}\exp(-\rho k)+\lambda_{\min}$ decays exponentially with $k$.
Step 4: Covering Number and Complexity Bound
Using the covering number argument and the fact that DRTR reduces effective neighborhood size from $d_{\text{original}}$ to $d_{\text{eff}}$, we obtain:
$\mathcal{R}_{n}(\mathcal{H}_{\text{DRTR}})\leq\mathcal{O}\left(\sqrt{\frac{d_{\text{eff}}\log n}{n}}\right),$ (22)
where the covering number $\mathcal{N}(\mathcal{H}_{\text{DRTR}},\epsilon)$ satisfies:
$\log\mathcal{N}(\mathcal{H}_{\text{DRTR}},\epsilon)\leq d_{\text{eff}}\cdot\log\left(\frac{L_{\text{DRTR}}B}{\epsilon}\right),$ (23)
and $B$ is the bound on node features.
Step 5: Final Bound via Standard Generalization Theory
Combining the Rademacher complexity bound with standard generalization theory (see, e.g., bartlett2002rademacher):
$\mathcal{L}_{\text{true}}\leq\mathcal{L}_{\text{emp}}+2\mathcal{R}_{n}(\mathcal{H}_{\text{DRTR}})+\mathcal{O}\left(\sqrt{\frac{\log(1/\delta)}{n}}\right).$ (24)
Substituting the complexity bound from Step 4 and the effective degree reduction from Step 2 completes the proof.
A.2 Auxiliary Lemmas for Generalization Bound
We first establish several key lemmas that will be used in the main proof. These lemmas control the error propagation through the multi-hop aggregation structure and establish the properties of our learning rate schedule.
Lemma A.3 (Properties of Learning Rate Weights).
For the learning rate $\alpha_{t}=\frac{H+1}{H+t}$ and the weights $\alpha_{i}^{t}$ defined as:
$\alpha_{0}^{t}=\prod_{j=1}^{t}(1-\alpha_{j}),\quad\alpha_{i}^{t}=\alpha_{i}\prod_{j=i+1}^{t}(1-\alpha_{j}),$ (25)
the following properties hold:
  1. (a)
    $\frac{1}{\sqrt{t}}\leq\sum_{i=1}^{t}\frac{\alpha_{i}^{t}}{\sqrt{i}}\leq\frac{2}{\sqrt{t}}$ for every $t\geq 1$.
  2. (b)
    $\max_{i\in[t]}\alpha_{i}^{t}\leq\frac{2H}{t}$ and $\sum_{i=1}^{t}(\alpha_{i}^{t})^{2}\leq\frac{2H}{t}$ for every $t\geq 1$.
  3. (c)
    $\sum_{t=i}^{\infty}\alpha_{i}^{t}=1+\frac{1}{H}$ for every $i\geq 1$.
Proof A.4 (Proof of Lemma A.0).
The proof follows from direct manipulation of the definition of $\alpha_{t}$ and $\alpha_{i}^{t}$.
Part (a): The proof is by induction on $t$. For $t=1$, we have $\sum_{i=1}^{1}\frac{\alpha_{i}^{1}}{\sqrt{i}}=\alpha_{1}=\frac{H+1}{H+1}=1$, so the statement holds. For $t\geq 2$, by the relationship $\alpha_{i}^{t}=(1-\alpha_{t})\alpha_{i}^{t-1}$ for $i=1,\ldots,t-1$, we have:
$\sum_{i=1}^{t}\frac{\alpha_{i}^{t}}{\sqrt{i}}=\frac{\alpha_{t}}{\sqrt{t}}+(1-\alpha_{t})\sum_{i=1}^{t-1}\frac{\alpha_{i}^{t-1}}{\sqrt{i}}.$ (26)
By induction and the fact that $\alpha_{t}=\frac{H+1}{H+t}$, the bounds follow.
Part (b): We have:
$\alpha_{i}^{t}=\frac{H+1}{i+H}\cdot\frac{i+1}{i+1+H}\cdots\frac{t-1}{t-1+H}\leq\frac{H+1}{i+H}\cdot\frac{t}{t+H}\leq\frac{2H}{t},$ (27)
where the last inequality uses $H\geq 1$. The second inequality follows since $\sum_{i=1}^{t}(\alpha_{i}^{t})^{2}\leq[\max_{i\in[t]}\alpha_{i}^{t}]\cdot\sum_{i=1}^{t}\alpha_{i}^{t}=[\max_{i\in[t]}\alpha_{i}^{t}]\cdot 1$.
Part (c): Using the identity for telescoping products:
$\sum_{t=i}^{\infty}\alpha_{i}^{t}=\frac{H+1}{i+H}\cdot\left(1+\frac{i}{i+1+H}+\frac{i}{i+1+H}\cdot\frac{i+1}{i+2+H}+\cdots\right)=\frac{H+1}{H}=1+\frac{1}{H}.$ (28)
This completes the proof.
Lemma A.5 (Recursion on Representation Error).
For any node $v\in\mathcal{V}$ and hop $k\in[K]$, let $t=N_{k}(v)$ and suppose node $v$ was processed at hop $k$ in iterations $t_{1},\ldots,t_{t}<t$. Then the representation error satisfies:
$(\mathbf{z}_{v}^{t}-\mathbf{z}_{v}^{*})=\sum_{i=1}^{t}\alpha_{i}^{t}\cdot(\mathbf{z}_{u_{i}}^{t_{i}}-\mathbf{z}_{u_{i}}^{*})+\beta_{t},$ (29)
where $\alpha_{i}^{t}$ are the learning rate weights, $u_{i}$ are the neighbors used in aggregation, and $\beta_{t}$ is a confidence bonus term.
Proof A.6 (Proof of Lemma A.1).
The proof follows from the update equation (4.1) and the definition of $\alpha_{i}^{t}$ in (4.2). By expanding the recursive update:
$\mathbf{z}_{v}^{t}=\alpha_{0}^{t}\mathbf{z}_{v}^{0}+\sum_{i=1}^{t}\alpha_{i}^{t}\left[\mathbf{W}^{(k)}\mathbf{x}_{u_{i}}+\mathbf{b}_{i}\right],$ (30)
where $\mathbf{b}_{i}$ is the confidence bonus. Subtracting the optimal representation $\mathbf{z}_{v}^{*}$ and using the Bellman-like optimality condition completes the proof.
Lemma A.7 (Bound on Representation Error).
There exists an absolute constant $c>0$ such that, for any $p\in(0,1)$, with probability at least 1p1-p, the following holds simultaneously for all $(v,k,t)\in\mathcal{V}\times[K]\times[T]$:
$0\leq(\mathbf{z}_{v}^{t}-\mathbf{z}_{v}^{*})^{\top}\mathbf{z}_{v}^{*}\leq\alpha_{0}^{t}H+\sum_{i=1}^{t}\alpha_{i}^{t}(\mathbf{z}_{u_{i}}^{t_{i}}-\mathbf{z}_{u_{i}}^{*})^{\top}\mathbf{z}_{u_{i}}^{*}+\beta_{t},$ (31)
where $\beta_{t}=2\sum_{i=1}^{t}\alpha_{i}^{t}b_{i}\leq 4c\sqrt{H^{3}\iota/t}$ and $\iota=\log(SAT/p)$.
Proof A.8 (Proof of Lemma A.2).
The proof uses Azuma-Hoeffding concentration inequality and the fact that the martingale difference sequence has bounded variance. For each fixed $(v,k)$, the sequence $\{(\mathbf{z}_{v}^{t}-\mathbf{z}_{v}^{*})^{\top}\mathbf{z}_{v}^{*}\}_{t=1}^{T}$ forms a martingale with respect to the filtration generated by previous updates. By Azuma-Hoeffding:
$\mathbb{P}\left[\sum_{i=1}^{t}\alpha_{i}^{t}\xi_{i}>c\sqrt{H^{3}\iota/t}\right]\leq\frac{p}{SAT},$ (32)
where $\xi_{i}$ are the martingale differences. A union bound over all $(v,k,t)$ completes the proof.
A.3 Graph Perturbation and Error Decomposition
For ground-truth graph $\mathcal{G}^{*}$ with structural smoothness ($\|\mathbf{x}_{u}-\mathbf{x}_{v}\|_{2}^{2}\leq\delta$ for $(u,v)\in\mathcal{E}^{*}$) and observed graph $\mathcal{G}$ with $\epsilon$-fraction noisy edges, the representation error decomposes as:
$\|\mathbf{z}_{v}-\mathbf{z}_{v}^{*}\|_{2}\leq\underbrace{\|\mathbf{z}_{v}-\mathbf{z}_{v}^{\text{DRTR}}\|_{2}}_{\text{Noise attenuation}}+\underbrace{\|\mathbf{z}_{v}^{\text{DRTR}}-\mathbf{z}_{v}^{*}\|_{2}}_{\text{Bias due to pruning}}.$ (33)
A.4 Bounded Noise Reduction via Distance Thresholding
Proposition A.9 (Noise Attenuation).
For $\ell$-Lipschitz features over clean graph $\mathcal{G}^{*}$, pruning edges with $\|\mathbf{x}_{u}-\mathbf{x}_{v}\|_{2}^{2}>\alpha$ yields:
$\mathrm{Var}[\mathbf{z}_{v}^{\text{DRTR}}]\leq\mathrm{Var}[\mathbf{z}_{v}]-\epsilon\cdot C(\alpha,\ell),$ (34)
where $C(\alpha,\ell)=\min_{(u,v)\in\mathcal{E}_{\text{noisy}}}\alpha_{vu}^{2}\cdot\mathrm{Var}[\mathbf{W}\mathbf{x}_{u}]>0$ increases with $\alpha$.
A.5 Optimization Analysis and Convergence Guarantees
A.5.1 Objective Function Decomposition
The DRTR optimization problem:
$\min_{\theta}\mathcal{L}(\theta)=\mathcal{L}_{\text{classification}}(\theta)+\lambda_{1}\mathcal{L}_{\text{DR}}(\theta)+\lambda_{2}\mathcal{L}_{\text{TR}}(\theta)+\lambda_{3}\mathcal{L}_{\text{regularization}}(\theta),$ (35)
where:
$\mathcal{L}_{\text{classification}}(\theta)$ $=-\sum_{v\in\mathcal{V}_{L}}y_{v}\log\hat{y}_{v}$ (36)
$\mathcal{L}_{\text{DR}}(\theta)$ $=\sum_{k=1}^{K}\sum_{v\in\mathcal{V}}\sum_{u\in\mathcal{N}^{(k)}(v)}\max(0,d_{vu}^{(k)}-\alpha_{k})$ (37)
$\mathcal{L}_{\text{TR}}(\theta)$ $=\sum_{(v,u)\in\mathcal{E}_{\text{candidate}}}\max(0,\beta-\text{sim}_{vu})\cdot P(\text{add edge }(v,u))$ (38)
$\mathcal{L}_{\text{regularization}}(\theta)$ $=\sum_{k=1}^{K}\|\mathbf{W}^{(k)}\|_{F}^{2}+\sum_{k=1}^{K}\phi_{k}^{2}$ (39)
A.5.2 Gradient Analysis
The gradient:
$\nabla\mathcal{L}(\theta)=\nabla\mathcal{L}_{\text{classification}}(\theta)+\lambda_{1}\nabla\mathcal{L}_{\text{DR}}(\theta)+\lambda_{2}\nabla\mathcal{L}_{\text{TR}}(\theta)+\lambda_{3}\nabla\mathcal{L}_{\text{regularization}}(\theta),$ (40)
with component gradients:
$\frac{\partial\mathcal{L}_{\text{DR}}}{\partial\mathbf{W}^{(k)}}$ $=\sum_{v\in\mathcal{V}}\sum_{u\in\mathcal{N}^{(k)}(v)}\mathbb{I}[d_{vu}^{(k)}>\alpha_{k}]\cdot\frac{\partial d_{vu}^{(k)}}{\partial\mathbf{W}^{(k)}},$ (41)
$\frac{\partial\mathcal{L}_{\text{TR}}}{\partial\omega_{i}}$ $=\sum_{(v,u)\in\mathcal{E}_{\text{candidate}}}\mathbb{I}[\beta-\text{sim}_{vu}>0]\cdot P(\text{add edge }(v,u))\cdot\frac{\partial\text{sim}_{vu}}{\partial\omega_{i}}.$ (42)
A.5.3 Convergence Rate Analysis
Theorem A.10 (Convergence Rate of DRTR).
Under the assumptions that the objective function $\mathcal{L}(\theta)$ is $L$-smooth (i.e., $\|\nabla\mathcal{L}(\theta_{1})-\nabla\mathcal{L}(\theta_{2})\|_{2}\leq L\|\theta_{1}-\theta_{2}\|_{2}$), the gradients are bounded such that $\mathbb{E}[\|\nabla\mathcal{L}(\theta)\|_{2}^{2}]\leq G^{2}$, and the learning rate satisfies $\eta_{t}=\frac{\eta_{0}}{\sqrt{t}}$ for some $\eta_{0}>0$, the DRTR algorithm achieves:
$\min_{t=1,\ldots,T}\mathbb{E}[\|\nabla\mathcal{L}(\theta_{t})\|_{2}^{2}]\leq\frac{2(\mathcal{L}(\theta_{0})-\mathcal{L}^{*})}{\eta_{0}\sqrt{T}}+\frac{\eta_{0}G^{2}}{2\sqrt{T}},$
where $\mathcal{L}^{*}$ is the optimal objective value.
Proof A.11 (Convergence Proof).
Since DRTR preserves $L$-smoothness and bounded gradients, standard SGD analysis applies. The proof proceeds in several steps.
Step 1: Descent Lemma By $L$-smoothness of $\mathcal{L}$:
$\mathcal{L}(\theta_{t+1})\leq\mathcal{L}(\theta_{t})+\langle\nabla\mathcal{L}(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle+\frac{L}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}.$ (43)
Step 2: Update Rule The DRTR update rule is:
$\theta_{t+1}=\theta_{t}-\eta_{t}\nabla\mathcal{L}(\theta_{t}).$ (44)
Step 3: Expected Descent Taking expectation over the stochastic gradient:
$\mathbb{E}[\mathcal{L}(\theta_{t+1})]\leq\mathcal{L}(\theta_{t})-\eta_{t}\|\nabla\mathcal{L}(\theta_{t})\|_{2}^{2}+\frac{L\eta_{t}^{2}G^{2}}{2},$ (45)
where $G^{2}=\mathbb{E}[\|\nabla\mathcal{L}(\theta_{t})\|_{2}^{2}]$ is the bound on gradient variance.
Step 4: Telescoping Sum Summing over $t=1,\ldots,T$ and rearranging:
$\sum_{t=1}^{T}\eta_{t}\|\nabla\mathcal{L}(\theta_{t})\|_{2}^{2}\leq\mathcal{L}(\theta_{0})-\mathcal{L}^{*}+\frac{LG^{2}}{2}\sum_{t=1}^{T}\eta_{t}^{2},$ (46)
where $\mathcal{L}^{*}$ is the optimal objective value.
Step 5: Learning Rate Analysis With $\eta_{t}=\frac{\eta_{0}}{\sqrt{t}}$:
$\sum_{t=1}^{T}\eta_{t}=\eta_{0}\sum_{t=1}^{T}\frac{1}{\sqrt{t}}\geq\eta_{0}\sqrt{T},$ (47)
and
$\sum_{t=1}^{T}\eta_{t}^{2}=\eta_{0}^{2}\sum_{t=1}^{T}\frac{1}{t}\leq\eta_{0}^{2}\log T.$ (48)
Step 6: Final Bound Combining the above results:
$\min_{t=1,\ldots,T}\mathbb{E}[\|\nabla\mathcal{L}(\theta_{t})\|_{2}^{2}]\leq\frac{\mathcal{L}(\theta_{0})-\mathcal{L}^{*}+\frac{LG^{2}\eta_{0}^{2}\log T}{2}}{\eta_{0}\sqrt{T}}=\mathcal{O}\left(\frac{1}{\sqrt{T}}\right).$ (49)
This completes the convergence rate analysis.
A.6 Computational Optimization Analysis
A.6.1 Time Complexity Optimization
The computational complexity of DRTR is optimized through component-wise analysis. The Distance Recomputator (DR) module performs distance computation:
$T_{\text{DR}}=K\cdot\sum_{v\in\mathcal{V}}\sum_{u\in\mathcal{N}^{(k)}(v)}\mathcal{O}(d)=\mathcal{O}(K\cdot|\mathcal{E}|\cdot d),$ (50)
with thresholding operations:
$T_{\text{threshold}}=K\cdot|\mathcal{V}|\cdot\mathcal{O}(\log|\mathcal{V}|)=\mathcal{O}(K\cdot|\mathcal{V}|\cdot\log|\mathcal{V}|),$ (51)
yielding total DR complexity $T_{\text{DR}}^{\text{total}}=\mathcal{O}(K\cdot|\mathcal{E}|\cdot d)$ (thresholding is dominated by distance computation).
The Topology Reconstructor (TR) module uses approximate $k$-nearest neighbor search, reducing candidate edges from $\mathcal{O}(|\mathcal{V}|^{2})$ to $\mathcal{O}(|\mathcal{V}|\cdot k)$:
$T_{\text{TR}}=\mathcal{O}(|\mathcal{V}|\cdot k\cdot\log|\mathcal{V}|)+\mathcal{O}(|\mathcal{V}|\cdot k\cdot d)=\mathcal{O}(|\mathcal{V}|\cdot k\cdot d),$ (52)
where $k\ll|\mathcal{V}|$ (typically $k=50$).
The multi-hop aggregation component involves attention computation:
$T_{\text{attention}}=K\cdot\sum_{v\in\mathcal{V}}\sum_{u\in\mathcal{N}^{(k)}(v)}\mathcal{O}(d^{2})=\mathcal{O}(K\cdot|\mathcal{E}|\cdot d^{2}),$ (53)
and message passing:
$T_{\text{message}}=K\cdot|\mathcal{V}|\cdot\mathcal{O}(d^{2})=\mathcal{O}(K\cdot|\mathcal{V}|\cdot d^{2}).$ (54)
Combining these components, the overall time complexity is:
$T_{\text{DRTR}}=T_{\text{attention}}+T_{\text{message}}+T_{\text{DR}}^{\text{total}}+T_{\text{TR}}=\mathcal{O}(K\cdot|\mathcal{E}|\cdot d^{2}+|\mathcal{V}|\cdot k\cdot d),$ (55)
which scales linearly with graph size $|\mathcal{V}|$ and edge count $|\mathcal{E}|$, making DRTR suitable for large-scale graphs.
A.6.2 Computational Speedup Analysis
Compared to naive multi-hop aggregation without optimization, DRTR achieves computational speedup through:
$\text{Speedup}=\frac{T_{\text{naive}}}{T_{\text{DRTR}}}=\frac{\mathcal{O}(K\cdot|\mathcal{V}|^{2}\cdot d^{2})}{\mathcal{O}(K\cdot|\mathcal{E}|\cdot d^{2}+|\mathcal{V}|\cdot k\cdot d)}\approx\frac{|\mathcal{V}|^{2}}{|\mathcal{E}|+|\mathcal{V}|\cdot k/d},$ (56)
where $|\mathcal{E}|\ll|\mathcal{V}|^{2}$ for sparse graphs, yielding significant speedup. The effective neighborhood reduction factor is:
$\rho=\frac{d_{\text{eff}}}{d_{\text{original}}}=\frac{\sum_{v}|\mathcal{N}_{\text{eff}}(v)|}{\sum_{v}|\mathcal{N}_{\text{original}}(v)|}\ll 1,$ (57)
leading to computational reduction proportional to $\rho$.
A.6.3 Space Complexity
The space complexity is:
$S_{\text{DRTR}}=\mathcal{O}(|\mathcal{V}|\cdot d)+\mathcal{O}(K\cdot|\mathcal{E}|)+\mathcal{O}(K\cdot|\mathcal{E}|)=\mathcal{O}(|\mathcal{V}|\cdot d+K\cdot|\mathcal{E}|),$ (58)
dominated by feature matrices and sparse attention/distance matrices.
A.7 Stability Analysis
Theorem A.12 (Stability of DRTR).
Let $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ be two graphs differing in at most $\Delta$ edges. Let $\mathbf{Z}_{1}$ and $\mathbf{Z}_{2}$ be the representations learned by DRTR on $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, respectively. Then:
$\|\mathbf{Z}_{1}-\mathbf{Z}_{2}\|_{F}\leq C\cdot\Delta\cdot\sqrt{|\mathcal{V}|},$
where $C$ is a constant depending on the Lipschitz constant and graph properties.
Proof A.13 (Stability Proof).
By Lipschitz continuity and bounded edge influence, with $|\mathcal{E}_{1}\triangle\mathcal{E}_{2}|\leq\Delta$:
$\|\mathbf{Z}_{1}-\mathbf{Z}_{2}\|_{F}\leq\sum_{k=1}^{K}\gamma_{k}\|\mathbf{H}_{1}^{(k)}-\mathbf{H}_{2}^{(k)}\|_{F}\leq\sum_{k=1}^{K}\gamma_{k}L_{k}\cdot\Delta\cdot\sqrt{|\mathcal{V}|}=C\cdot\Delta\cdot\sqrt{|\mathcal{V}|},$
where $C=\sum_{k=1}^{K}\gamma_{k}L_{k}$ and $L_{k}$ is the Lipschitz constant for hop $k$.
A.8 Additional Optimization Properties
Proposition A.14 (Convexity).
$\mathcal{L}_{\text{regularization}}$ is strongly convex, $\mathcal{L}_{\text{DR}}$ is convex in distance parameters, while the overall objective is non-convex due to classification loss.
Proposition A.15 (Saddle Point Avoidance).
Under mild conditions, DRTR avoids saddle points with high probability due to non-zero gradients and stochastic optimization.
A.9 Hyperparameter Optimization
The learning rate schedule:
$\eta_{t}=\frac{\eta_{0}}{\sqrt{t}}\cdot\exp(-\mu t),$ (59)
where $\eta_{0}$ is initial rate and $\mu$ controls decay. Regularization parameters satisfy:
$\lambda_{1}+\lambda_{2}+\lambda_{3}=1,\quad\lambda_{i}\geq 0.$ (60)
A.10 Extension to Other Graph Learning Tasks
DRTR extends to link prediction and graph classification by adapting similarity computation and pooling operations. For link prediction, edge probabilities are computed from learned representations using the topology reconstruction similarity function. For graph classification, node representations are aggregated via graph-level pooling (mean/attention/hierarchical), with distance recomputation adapted for inter-graph similarities.
A.11 Implementation Details
A.11.1 Layer Normalization Details
Layer normalization is applied after each hop aggregation to ensure numerical stability and prevent vanishing gradients. The detailed formulation is:
$\tilde{\mathbf{h}}_{v}^{(k)}=\text{LayerNorm}\left(\mathbf{h}_{v}^{(k)}\right)=\frac{\mathbf{h}_{v}^{(k)}-\mu_{v}^{(k)}}{\sigma_{v}^{(k)}+\epsilon},$ (61)
where $\mu_{v}^{(k)}=\frac{1}{d}\sum_{i=1}^{d}h_{v,i}^{(k)}$ is the mean, $\sigma_{v}^{(k)}=\sqrt{\frac{1}{d}\sum_{i=1}^{d}(h_{v,i}^{(k)}-\mu_{v}^{(k)})^{2}}$ is the standard deviation, and $\epsilon=10^{-5}$ is a small constant to prevent division by zero.
A.11.2 Numerical Stability
We implement gradient clipping ($\|\nabla\mathcal{L}\|_{2}\leq\tau_{\text{clip}}$), layer normalization (Eq. (4)), and temperature scaling ($\tau_{k}=\tau_{0}\exp(-\eta k)$) to ensure numerical stability during training.
A.11.3 Computational Optimization
Sparse matrix operations reduce memory from $\mathcal{O}(|\mathcal{V}|^{2})$ to $\mathcal{O}(|\mathcal{E}|)$. Batch processing optimizes distance computation:
$T_{\text{batch}}=\left\lceil\frac{|\mathcal{V}|}{B}\right\rceil\cdot\mathcal{O}(B\cdot k\cdot d)=\mathcal{O}(|\mathcal{V}|\cdot k\cdot d),$ (62)
where $B$ is batch size. Approximate $k$-NN search reduces TR complexity from $\mathcal{O}(|\mathcal{V}|^{2}\cdot d)$ to $\mathcal{O}(|\mathcal{V}|\cdot k\cdot d)$.
A.11.4 Implementation Settings
Table LABEL:tab:implementation-settings summarizes the key hyperparameters and implementation settings used in our DRTR framework across different datasets.
[…]
tab:implementation-settings Setting Category Parameter Value Training Learning Rate 0.005 Weight Decay 0.0005 Epochs 1000 Early Stopping Patience 100 Batch Size Full graph Numerical Stability Gradient Clipping Threshold 1.0 Layer Normalization $\epsilon$ $10^{-5}$ Initial Temperature $\tau_{0}$ 1.0 Temperature Decay Rate $\eta$ 0.1 Distance Recomputator Initial Scaling $\lambda_{0}$ 0.1 Scaling Decay Rate $\rho$ 0.05 Minimum Scaling $\lambda_{\min}$ 0.01 Pruning Percentile $p$ 0.75 Topology Reconstructor Similarity Threshold $\beta$ 0.5 Edge Addition Probability $\theta$ 0.6 Approximate NN Search (k) 50 Multi-Hop Aggregation Maximum Hops $K$ 3 Hidden Dimension $d^{\prime}$ 64 Attention Head Dimension 8 Computation Sparse Matrix Format CSR Distance Batch Size 1024
Table 5: Implementation settings and hyperparameters for DRTR framework.